OpenStudy (anonymous):

Restriction help!! A box with a square base and no top must have a volume of 10000cm^3. Determine the dimensions of the box that minimize the amount of materials used. The smallest dimension possible is 5cm What is the restriction in here? I already found the answer h = 13.6 and w=l = 27.1 but dont know what is the restrictions

OpenStudy (saifoo.khan):

@mathteacher1729

OpenStudy (anonymous):

restriction is \(w^2h=3000\)

OpenStudy (anonymous):

where is that 3000 from?? first I thought is 100 because 100x1000 gives 10000 (so cant have any volume) but when i plug it into the surface area equation, it gives a huge number..

OpenStudy (anonymous):

oh because i wasn't paying attention doh should have been \(w^2h=1000\)

OpenStudy (anonymous):

can u tell me how to get it...kinda confuse

OpenStudy (anonymous):

you are given that the volume is fixed, that is it must be 1000 square whatever and the base is a square with area \(w^2\) assuming you are using "w" as the variable representing the lenght of the base, and the height is "h" if you use that variable for height volume is therefore \(w^2h\) area of the base times the height.

OpenStudy (mathteacher1729):

The very first problem on this page is almost identical to your original problem. :) http://www.cliffsnotes.com/study_guide/Maximum-Minimum-Problems.topicArticleId-39909,articleId-39895.html

OpenStudy (anonymous):

The equation I have is S = x^2 + 40000/x (when i combine the 2 equations together) so if i plug in x = 1000 , then i get 1000040 for surface area so doesnt seem restricted?

OpenStudy (anonymous):

oh really, i will take a look at it! thks mathteacher

OpenStudy (anonymous):

so you have two representations for the volume, one in terms of variables \(w^2h\) and the other a number you know, namely 10000, giving \[w^2h=1000\]

OpenStudy (anonymous):

isnt it w^2h = 10000

OpenStudy (anonymous):

@mathteacher1729 but that question only shows the min value, but not the restrictions

OpenStudy (anonymous):

restriction is first line \[x^2h=100\]

OpenStudy (anonymous):

or rather \[x^2h=1000\]